Circular No. 9033

Central Bureau for Astronomical Telegrams INTERNATIONAL ASTRONOMICAL UNION

Mailstop 18, Smithsonian Astrophysical Observatory, Cambridge, MA 02138, U.S.A.

IAUSUBS@CFA.HARVARD.EDU or FAX 617-495-7231 (subscriptions) CBAT@CFA.HARVARD.EDU (science)

        URL http://www.cfa.harvard.edu/iau/cbat.html ISSN 0081-0304

Phone 617-495-7440/7244/7444 (for emergency use only)

COMET C/2009 F5 (McNAUGHT) R. H. McNaught reports his discovery of another comet on CCD sur-

vey images obtained with the 0.5-m Uppsala Schmidt reflector at Siding Spring (discovery observation tabulated below), the object showing a 0′.6 coma that appears extended toward the west, though the full extent and existence of a tail is diffult to judge due to the crowded star field. Following posting on the Minor Planet Center’s ‘NEOCP’ webpage, other CCD astrometrists have also commented on the object’s cometary appearance. J. E. McGaha (Tucson, AZ, U.S.A., 0.36-m f/10 Schmidt-Cassegrain reflector, Mar. 21.4 UT) relates that five stacked 60-s exposures show a bright 6′′ nuclear condensation surrounded by a faint round 16′′ coma and no tail. Three stacked 60-s images taken with the Mt. John 1.0-m f/7.7 reflector by A. C. Gilmore on Mar. 21.5 (measured by P. M. Kilmartin) show a circular coma of diameter 13′′ with a stellar condensation at its center but no tail; however, three stacked 120-s images taken on Mar. 22.4 at Mt. John show a circular 20′′ coma with the central starlike condensation and a possible faint, broad tail 90′′ long in p.a. ∼ 230^{o}. E. Guido (Castellammare di Stabia, Italy) writes that twenty co-added, unfiltered 120-s CCD exposures taken remotely with 0.25-m f/3.4 reflector near Mayhill, NM, U.S.A., by G. Sostero, P. Camilleri, and himself on Mar. 22.4 reveal a coma diameter of ∼ 30′′ and a strong central condensation. J. W. Young (Table Mountain 0.61-m f/16 Cassegrain reflector, Mar. 22.4) reports that his exposures taken at low altitude show a fuzzy appearance with a size around 10′′–12′′. R. Ligustri (Talmassons, Udine, Italy, remotely using a 0.35-m f/7 reflector at Grove Creek Observatory, Trunkey, N.S.W., Australia, Mar. 22.63–22.66) finds a coma of diameter ∼ 15′′, elongated toward p.a. 250^{o}, on ten co-added 60-s exposures.

2009 UT  α_{2000}  δ_{2000}  Mag. Observer
Mar. 20.65563 15^{h}14^{m}43.^{s}77 −36^{o}41′46′′.5 15.9 McNaught

The available astrometry, the following very preliminary parabolic orbital elements by B. G. Marsden, and an ephemeris appear on MPEC 2009-F60.

T = 2008 Nov. 8.164 TT  ω = 299.^{o}227
Ω = 219.027 }2000.0
q = 2.29052 AU  i = 85.783
2009 March 22  \begin{matrix}c \\ ◯ \end{matrix} Copyright 2009 CBAT  Daniel W. E. Green